OpenStudy (anonymous):

A square with sides of 3sqrt2 is inscribed in a circle. What is the area of one of the sectors formed by the radii to the vertices of the square?

OpenStudy (anonymous):

0.75*Pi=0.75*3.14159=2.355 (aprox)

OpenStudy (anonymous):

3pi/4